# MMAR: Towards Lossless Multi-Modal Auto-Regressive Probabilistic Modeling

### v2

@@ 3.2 Optimal Diffusion Model Parameterization for Low-Precision Training
In the era of large language models, training with low precision data type, such as bfloat16, has become increasingly popular. However, the training and inference process of a diffusion model is relatively sensitive to numerical precision. Moreover, in an auto-regressive framework, the image tokens are sampled sequentially, requiring even more precise sampling for each image token to reduce the overall error accumulation. Therefore, handling the numerical error in the diffusion process modeling should be emphasized when integrating diffusion loss into LLMs.
From the example below, we can clearly illustrate the effect of the numerical error in diffusion models. In eq.3, the diffusion model is parameterized as $\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})$, known as “$\epsilon$-prediction”, predicting the noise $\epsilon$ that is added to the data. Floating-point representation causes numerical error’s amplitude proportional to a number’s magnitude. This can be modeled by scaling the prediction by a factor of $(1+\delta)$, where $\delta$ is the relative error. In this way, we can write DDIM (Song et al., 2021) sampling with numerical error as follows:
$\tilde{x}^{(t-1)}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x^{(t)}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta).$ (5)
@@
$\tilde{x}^{(t-1)}=x^{(t-1)}+(\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}})\epsilon_{\theta}(x^{(t)},t,z_{i})\delta,$ (6)
where the first term $x^{(t-1)}$ is the ideal DDIM sampling term, and the second term is the numerical error term. Assuming $\epsilon_{\theta}(x^{(t)},t,z_{i})$ follows a standard normal distribution, the standard deviation of numerical error term can be calculated as $|\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}}|\delta$. When the signal-to-noise ratio (SNR) is high, i.e. $\bar{\alpha}_{t},\bar{\alpha}_{t-1}\to 1$, the numerical error has almost zero amplitude. However, when SNR is extremely low, i.e. $\bar{\alpha}_{t}=0$, and $\bar{\alpha}_{t-1}>0$, this term will explode to infinity, causing extreme numerical error.
Our goal is to minimize the numerical error term. To achieve this, we analyze all the factors that can determine the numerical error in Appendix A.2, and find that the only solution is to parameterize the diffusion model properly. By solving the numerical error minimization problem, we conclude that the v-prediction parameterization (Salimans & Ho, 2022) is the desired optimal parameterization method. Note that v-prediction is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models. To the best of our knowledge, our work is the first to derive v-prediction parameterization from the first principle of minimizing the numerical error in diffusion models. For more details, see Appendix A.2. Under v-prediction parameterization, the model predicts a linear combination of data $x_{i}$ and noise $\epsilon$:
$v^{(t)}_{i}=\sqrt{\bar{\alpha}_{t}}\mathbf{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}x_{i}$ (7)
@@
$L(x_{i})=\mathbb{E}_{x_{i},\epsilon,t}[||v_{i}^{(t)}-v_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}].$ (8)
3.3 Model Architecture and Training Strategy
3.3.1 Pipeline
To jointly model the probabilities of images and text at the token level, we employ an image tokenizer to extract image tokens and randomly mask some of them for image generation training, as depicted in fig. 2. Given the substantial distribution gap between images and text, directly scaling image tokens to match the LLM’s dimensions would pose a challenge to joint image-text probability modeling. To address this, we introduce a Latent Encoding process utilizing an EmbeddingViT module for the unmasked image tokens, which simplifies the modeling task. To facilitate the auto-regressive training of image tokens, we concatenate the known image tokens with the masked tokens and append the position embedding corresponding to their original positions. These tokens, after being processed by the VisProjector, are concatenated with the text tokens to generate the final image-text sequence. Then, an LLM is used to process the image-text sequence, whose output $z$ integrates information from both the preceding text and the unmasked image tokens. For image output, $z$ acts as a conditional input for the Diffusion MLP, which predicts the $v$ values of the noisy masked image tokens. The diffusion loss, as depicted in eq.8, is applied, enabling us to model the probability of the continuous image tokens. For text output, a linear lm-head processes $z$ to predict the next token’s logits, and the corresponding cross entropy loss is applied, allowing us to model the probability of the text tokens.
During image generation, to ensure consistency with the training scenario, we initialize all image tokens as mask tokens and randomly generate a set of image position sequences. Following a left-to-right order, we extract N condition tokens $z$ at each step and input them into the Diffusion MLP to generate the corresponding image tokens. Next, we assign these image tokens back to their respective positions in the original image token sequence, iterating until the entire image token set is obtained. Ultimately, we decode the image tokens into images using the image tokenizer, yielding the generated images.
3.3.2 Model Architecture
Tokenizer
MMAR employs the publicly available KL-16 image tokenizer from LDM (Rombach et al., 2022). This tokenizer processes an 256 $\times$ 256 image into 256 image tokens, each with 16 channels, and it is kept frozen during training.
EmbeddingViT
To achieve a more profound encoding of image tokens, we introduce the EmbeddingViT module, a ViT(Dosovitskiy et al., 2021) encoder with 16 layers and 1024 hidden state channels, processing image tokens into visual embeddings with stronger context awareness. Moreover, as the input consists of randomly shuffled image tokens, we integrate a learnable position embedding for each image token in EmbeddingViT, corresponding to its original position.
LLM
We initialize our LLM using parameters from the open-source QWen2 series models (Yang et al., 2024). To preserve text modeling capabilities, we keep the LLM parameters fixed during training and add PLoRA (Dong et al., 2024b) as the image expert, where only the image tokens in the input pass through the introduced LoRA (Hu et al., 2022) adapters. The PLoRA is applied for each linear layer within the original LLM. Considering the time cost, our ablation study employs QWen2-0.5B-Instruct, with $r=512$ for PLoRA. Furthermore, we use QWen2-7B-Instruct to explore our method’s scale-up capability with $r=1280$ for PLoRA. The corresponding MMAR models are denoted by MMAR-0.5B and MMAR-7B, respectively.
In the LLM, we adopt a bidirectional attention mechanism to enhance interaction among image tokens, rendering all image tokens mutually visible, as depicted in fig. 2. Meanwhile, the text portion retains causal attention. Additionally, to prevent interference with the auto-regressive training of image tokens in random order, we assign the same ROPE position id to all image tokens within all LLM layers. This strategy not only ensures the random order regression of image tokens but also mitigates the issue of text concentrating on closer image tokens in long text scenarios.
Diffusion MLP
Inspired by MAR, we employ a simple MLP to predict $v^{(t)}$. This MLP consists of a series of residual blocks, each comprising AdaLN (Peebles & Xie, 2023), a linear layer, SiLU, and another linear layer. The condition vector $z$ is added to the diffusion time embedding and incorporated through AdaLN. MMAR-0.5B utilizes a Diffusion MLP with 8 residual blocks and 1024 channels, while MMAR-7B employs a larger variant with 12 residual blocks and 2048 channels.
3.3.3 Training Strategy
Multi-Task Training
To accomplish joint image-text modeling, we simultaneously conducted text-to-image, image-to-text, and unconditional image generation tasks during training. In the image-to-text task, no mask tokens are assigned to the image part, allowing us to model $P(T|I)$ using the complete image tokens. Furthermore, the allocation ratio of text-to-image and unconditional image generation tasks is set to 9:1, facilitating efficient use of the cfg technique at the inference stage. To maintain a balance between tasks, we intuitively set the sample allocation ratio of image generation tasks and image understanding tasks to 1:1.
Two Stage Training
@@ 3.3.3 Training Strategy
To achieve a balance between image generation and understanding capabilities, our training process is carried out in two stages. The first stage utilizes large-scale, mid-quality data (as illustrated in fig. 2) to enhance the diversity of the model’s data distribution (Hu et al., 2024). By initially modeling this diverse data, we strengthen the model’s understanding capabilities. In the second stage, we employ a smaller volume of high-quality data to further improve the image generation capacity and refine the model’s comprehension of images.
We observe that when the training dataset is excessively large, preventing the model from iterating over each image hundreds of times, maintaining an image mask ratio within the range of [0.7, 1], as done in MAR, hinders the model’s ability to generate coherent image tokens. Consequently, in the second stage, we adjust the mask ratio to (0, 1]. The example is showed in Appendix A.5. The probability density curve of the two mask ratios are shown on the right side of fig. 2. Additionally, due to the small size of the diffusion mlp and the limited amount of high-quality data, we increased the number of timestep samples for the given $z$ to enhance learning efficiency. To further mitigate performance fluctuations in the model, we employed an exponential moving average (EMA) method with a momentum of 0.9999 in both stages.
3.3.4 Inference Strategy
Classifier-Free Guidance
During the inference stage, the model performs both text-based and non-text-based image generation tasks. The provided conditions are represented as $z_{c}$ and $z_{u}$, and the predicted $v$ is given by: $v=v_{\theta}(x^{(t)}|t,z_{u})+\omega*(v_{\theta}(x^{(t)}|t,z_{c})-v_{\theta}(x^{(t)}|t,z_{u}))$ (which has the same effect as $\epsilon=\epsilon_{u}+\omega*(\epsilon_{c}-\epsilon_{u})$ , with the mathematical derivation process detailed in the Appendix A.4), where $\omega$ denotes the guidance scale. Our method has been experimentally validated to support a large guidance scale, as compared to the noise prediction training approach. We hypothesize that this is due to the strong conditioning provided by $z$, which, coupled with the auto-regressive (AR) process, necessitates reducing accumulated error. Thus, each token must be generated as accurately as possible, a requirement that can be fulfilled with a large guidance scale.
Mixed-Precision Inference
To enhance numerical stability during image generation, particularly when using small step sizes in the DDPM sampling process, we perform the model’s forward pass in bfloat16 (matching the training precision) but cast the output to float32 before DDPM sampling. This mitigates potential rounding error without a significant computational overhead, therefore improving the sampling accuracy efficiently.
@@ 4.1 Dataset
We utilized the Capfusion-120M dataset for the image expert pretraining stage. This dataset is publicly accessible and comprises an extensive collection of web-based image-text pairs, designed to optimize noisy captions (Yu et al., 2023). In an effort to further improve the quality of the content generated, we executed a random sampling of 20M data points from the Capfusion dataset during our image expert fine-tuning stage. This was supplemented with a high-quality mixed dataset that included ImageNet-1K-1.2M (Deng et al., 2009), CC12M (Changpinyo et al., 2021), and LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap. We employ the open-source InternVL2-8B for recaptioning the CC12M and laion-aesthetics-12m datasets in English. Following LLaVA-v1.5 (Liu et al., 2023a), we use LLaVA-v1.5-mix-665K data for instruction tuning before each performance test for visual understanding.
@@ 4.2 Implementation Details
By default, we utilized the AdamW optimizer with betas (0.9, 0.95). The weight decay was consistently maintained in proportion to the learning rate. During the first training stage, a learning rate of 5e-5 was employed for a total of 4 epochs, with an initial warm-up phase of 0.5 epochs, followed by maintaining the learning rate at 5e-5. In the second stage, the 0.5B model maintained a 5e-5 learning rate and a 0.5 epoch warm-up, with training lasting for 3 epochs. For the larger 7B model, we initially applied a 2e-6 learning rate for 1.5 epochs, before transitioning to a 1e-6 learning rate for an additional 1.5 epochs. Furthermore, during the first stage, the total batch size for the smaller 0.5B model was 2496, while it was 1152 for the larger 7B model. In the second stage, the total batch size for the smaller 0.5B model was 768, and for the larger 7B model, it was 480. Notably, in both stages, only the PLoRA portion of the LLM parameters is trainable, and a consistent image resolution of 256x256 was used throughout. Our models utilize bfloat16 during training and inference. For DDPM sampling, we convert the model output to float32 for higher precision.
@@ 4.3 Comparison with Other Systems
Table 1: Comparison on visual understanding benchmarks. MMAR surpasses other joint image-text probabilistic models by a large margin, even with a small resolution of 256x256, approaching the performance of traditional MLLMs like LLaVA, which employ pretrained CLIP vision encoder.
Method LLM V-Token Res. MMB MMEP POPE SEED MM-Vet AVE@18Und.
LLaVA-1.5 Vicuna-1.5-7B CLIP 336 64.3 1510.7 85.9 58.6 31.1 47.08
EMU-2 LLaMA-13B CLIP 448 62.8 48.5
SEED-X LLaMA-13B CLIP dynamic 75.4 1435.7 84.2
DreamLLM Vicuna-7B CLIP 224 58.2 36.6
Chameleon-7B 7B from scratch vq-vae 512 13.32 125.8 30.86 34.61 7.34 18.34
Transfusion* Qwen-2-0.5B vae 256 29.47 594.3 66.90 42.40 13.90 28.26
Show-o Phi-1.5B CLIP 336 42.44 1182.7 84.50 51.61 20.87 33.06
VILA-U LLaMA-2-7B vq-vae 256 1336.2 83.9 56.3 27.7
MMAR-0.5B Qwen-2-0.5B vae 256 48.45 882.1 70.74 55.70 18.49 34.56
MMAR-7B Qwen-2-7B vae 256 66.32 1393.9 83.02 64.52 27.80 46.52
@@
Visual Understanding. As depicted in table 1, we thoroughly gauge MMAR’s performance in visual understanding, employing VLMEvalKit (Duan et al., 2024) to perform extensive evaluations on prevalent visual understanding benchmarks, encompassing a total of 18 such assessments (average score denoted by “AVE@18Und.” in table 1, for details see Appendix A.7.) including MMB (Liu et al., 2023b), MME (Fu et al., 2023), POPE (Li et al., 2023c), SEED (Li et al., 2023a), MM-Vet (Yu et al., 2024), among others. Our method outperforms other joint image-text probabilistic models by a large margin, including Chameleon-7B (Team, 2024), Show-o (Xie et al., 2024), VILA-U (Wu et al., 2024) and our re-implemented version of Transfusion (denoted by “Transfusion*”), approaching the performance of traditional MLLMs like LLaVA, which employ pretrained CLIP vision encoder. Even without using any pre-trained CLIP or diffusion models and with small resolution of 256 $\times$ 256, MMAR-7B presents comparable or even better performance when compared to methods using pre-trained clip and diffusion models, including EMU-2 (Sun et al., 2024c), SEED-X (Ge et al., 2024), and DreamLLM (Dong et al., 2024a).
Visual Generation. We showcase the zero-shot FID of MMAR, evaluated on the MSCOCO 30k dataset, in Table 2. Our model’s performance is discernibly on par with current robust generative models, a noteworthy achievement for MMAR given its minimal training epochs. Repeated exposure to identical data can substantially enhance a model’s generative quality, as exemplified by MAR, which trains for 400 epochs on the 1.2M ImageNet dataset, and Show-o, which undergoes approximately 49 epochs of training on the 35M dataset.

### v3

@@ 3.2 Optimal Diffusion Model Parameterization for Low-Precision Training
In the era of large language models, training with low precision data type, such as bfloat16, has become increasingly popular. However, the training and inference process of a diffusion model is relatively sensitive to numerical precision, we provide direct empirical demonstration in Appendix 6.8. Moreover, in an auto-regressive framework, the image tokens are sampled sequentially, requiring even more precise sampling for each image token to reduce the overall error accumulation. Therefore, handling the numerical error in the diffusion process modeling should be emphasized when integrating diffusion loss into LLMs.
From the example below, we can clearly illustrate the effect of the numerical error in diffusion models. In Eq. 3, the diffusion model is parameterized as $\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})$, known as “$\epsilon$-prediction”, predicting the noise $\epsilon$ that is added to the data. Floating-point representation causes numerical error’s amplitude proportional to a number’s magnitude. This can be modeled by scaling the prediction by a factor of $(1+\delta)$, where $\delta$ is the relative error. In this way, we can write DDIM [51] sampling with numerical error as follows:
$\tilde{x}^{(t-1)}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x^{(t)}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta)}{\sqrt{\bar{\alpha}_{t}}}\right)$
$+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta).$ (5)
@@
$\tilde{x}^{(t-1)}=$ $(\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}})\epsilon_{\theta}(x^{(t)},t,z_{i})\delta$
$+x^{(t-1)},$ (6)
where the first term is the numerical error term, and the second term $x^{(t-1)}$ is the ideal DDIM sampling term. Let $\sigma_{t}$ be the standard deviation of $\epsilon_{\theta}(x^{(t)},t,z_{i})$, then the standard deviation of numerical error term can be calculated as $|\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}}|\delta\sigma_{t}$. When the signal-to-noise ratio (SNR) is high, i.e. $\bar{\alpha}_{t},\bar{\alpha}_{t-1}\to 1$, the numerical error has almost zero amplitude. However, when SNR is extremely low, i.e. $\bar{\alpha}_{t}=0$, and $\bar{\alpha}_{t-1}>0$, this term will explode to infinity, causing extreme numerical error.
Our goal is to minimize the numerical error term. To achieve this, we analyze most of the factors that can determine the numerical error in Appendix 6.2, and find that an effective solution is to parameterize the diffusion model properly. By solving the numerical error minimization problem, we conclude that the v-prediction parameterization [49] is the desired optimal parameterization method. Note that v-prediction is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models. To the best of our knowledge, our work is the first to derive v-prediction parameterization from the first principle of minimizing the numerical error in diffusion models. For more details, see Appendix 6.2. Under v-prediction parameterization, the model predicts a linear combination of data $x_{i}$ and noise $\epsilon$:
$v^{(t)}_{i}=\sqrt{\bar{\alpha}_{t}}\mathbf{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}x_{i}.$ (7)
@@
$L_{i}=\mathbb{E}_{x_{i},\epsilon,t}[||v_{i}^{(t)}-v_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}].$ (8)
3.3 Implementation
3.3.1 Model Design
Overall, our model follows the auto-regressive (AR) modeling paradigm for both image and text. For the text part, we leverage a pre-trained decoder-only LLM [63] for the modeling of text probability $P(T)$. While for the image part, we pick the masked AR implementation [30], which has been demonstrated to perform much better than decoder-only transformer implementation, meanwhile maintaining the form of $p_{\theta}(x_{i}|x_{<i})$ modeling. Our design goal is to embed the masked AR into the pre-trained LLM to realize a unified $f_{\theta}(\cdot)$ for multi-modal input and output. Generally, a masked AR implementation includes a bi-directional encoder, which encodes the known image tokens $x_{<i}$, and a bi-directional decoder, which decodes the masked image token $x_{i}$ given $x_{<i}$ and the mask position embeddings. Our approach is shown in Fig. 3. Firstly, we introduce an EmbeddingViT module to serve as the role of the encoder. Secondly, we directly embed the decoder into the LLM. This is achieved by adding PLoRA-based [10] image expert into the LLM, and setting the attention of the image region to be bi-directional. We also set the same ROPE position ID for all image tokens to ensure that only the mask position embedding decides the image token position information. In this way, the decoder-only LLM behaves just like the masked AR decoder when the PLoRA is activated. Meanwhile, thanks to the pre-trained LLM, the decoding of image tokens can be conditioned on the preceding text. To calculate the image diffusion loss, a Diffusion MLP module is adopted. This module takes the embedding of the image part output by the LLM, denoted by $z_{i}=f_{\theta}(x_{<i})$, as a conditional input to predict $v_{i}^{(t)}$ according to Eq. 7, ultimately achieving the loss of Eq. 8. For detailed module specifics, please refer to Appendix 6.1.
During training, we simultaneously learn two tasks: image understanding and image generation. For image understanding task, the sequence is arranged as [img][text], and we perform text auto-regressive training under the condition of all image tokens, achieving $P(T|I)$. For conditional image generation task, the sequence is arranged as [text][img], and we perform image masked AR training under the condition of preceding text tokens, achieving $P(I|T)$. Specifically, the image tokens are randomly masked with a controlled mask ratio decided by the training strategy. For unconditional image generation task, the sequence is arranged as [img] only, and we perform image masked AR training without any preceding text conditions, achieving $P(I)$. This also allows the application of Classifier-Free Guidance (CFG) techniques. Combining $P(I|T)$, $P(T|I)$, $P(I)$, and $P(T)$, we achieve the modeling of $P(I,T)=P(I|T)P(T)=P(T|I)P(I)$, form both text to image and image to text direction.
3.3.2 Two Stage Training
@@
Image understanding tasks are often applied in various complex scenarios. Therefore, the training requires highly diverse data, including but not limited to blurred images, small-sized images, images with greatly differing aspect ratios, etc. However, image generation tasks, on the other hand, place higher demands on data clarity, resolution, and even aesthetic criteria is one of the ways to evaluate generative performance. To achieve a balance between image generation and understanding capabilities, our training process is carried out in two stages. The first stage, called Image Expert Pretraining stage, utilizes large-scale, mid-quality data (as illustrated in Fig. 3) to enhance the diversity of the model’s data distribution [23]. By initially modeling this diverse data, we strengthen the model’s understanding capabilities. In the second stage, called Image Expert Fine-tuning stage, we employ a smaller volume of high-quality data to further improve the image generation capacity and refine the model’s comprehension of images.
Surprisingly, we find that the images generated by the model at the end of the first training stage all have holes, as shown in Fig.  4. We control the step-by-step generation of images tokens and ultimately find that the appearance of holes always occurs within the last thirty percent of the generation sequence. And this is because the first stage uses a [0.7, 1.0] mask ratio, which in the context of fewer training epochs, results in the model still not adapting to the image completion scene task of the last thirty percent. To address this issue, we set the lower bound of the mask ratio to 0 in the second stage and adjust the variance of the mask ratio distribution to enable the model to learn image generation tasks in various scenarios as efficiently as possible. The probability density curve of the two mask ratios are shown on the right side of Fig. 3. Ultimately, this achieves a rapid improvement in the model’s generative capabilities within a limited number of training epochs.
Figure 4: Training stage 2 with (0,1] mask ratio range addresses the issue of generating holes of training stage 1.
Table 1: Comparison on visual understanding benchmarks. MMAR surpasses other joint image-text probabilistic models even with a small resolution of 256x256, approaching the performance of traditional MLLMs like LLaVA, which employ pre-trained CLIP vision encoder.
Type Method LLM V-Token Res. MMB MMEP POPE SEED MM-Vet AVE@18Und.
Und. Only LLaVA-1.5 [34] Vicuna-1.5-7B CLIP 336 64.3 1510.7 85.9 58.6 31.1 47.08
Und. & Gen. Non-prob. EMU-2 [55] LLaMA-13B CLIP 448 62.8 48.5
SEED-X [16] LLaMA-13B CLIP dynamic 75.4 1435.7 84.2
Joint Prob. Models Chameleon-7B [57] 7B from scratch vq-vae 512 13.32 125.8 30.86 34.61 7.34 18.34
Transfusion* Qwen-2-0.5B vae 256 29.47 594.3 66.90 42.40 13.90 28.26
Show-o [62] Phi-1.5B CLIP 336 42.44 1182.7 84.50 51.61 20.87 33.06
VILA-U [61] LLaMA-2-7B vq-vae 256 1336.2 83.9 56.3 27.7
EMU-3 [60] 8B from scratch vq-vae 512 58.5 85.2 68.2 37.2
MMAR-0.5B Qwen-2-0.5B vae 256 48.45 882.1 70.74 55.70 18.49 34.56
MMAR-7B Qwen-2-7B vae 256 70.45 1486.9 84.02 68.63 30.64 48.25
@@ 4.1 Dataset
We utilized the Capfusion-120M dataset [64] for the image expert pretraining stage. This dataset is publicly accessible and comprises an extensive collection of web-based image-text pairs, designed to optimize noisy captions. In an effort to further improve the quality of the content generated, we executed a random sampling of 20M data points from the Capfusion dataset during our image expert fine-tuning stage. This was supplemented with a high-quality mixed dataset that included CC12M [2], and LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap. We employ the open-source InternVL2-8B [7] for recaptioning the CC12M and laion-aesthetics-12m datasets in English. Following LLaVA-v1.5 [33], we use LLaVA-v1.5-mix-665K data for instruction tuning before each performance test for visual understanding.
@@ 4.2 Implementation Details
By default, we used the AdamW [37] optimizer with betas (0.9, 0.95) and weight decay proportional to the learning rate. In the first training stage, a learning rate of 5e-5 was used for 4 epochs (with a 0.5-epoch warm-up) for the 0.5B model, and 2e-5 for 3 epochs (with a 0.5-epoch warm-up) for the 7B model. The second stage used 5e-5 for 3 epochs (with a 0.5-epoch warm-up) for the 0.5B model, and 2e-5 for 3 epochs (with a 0.5-epoch warm-up) for the 7B model. Batch sizes were 2496 for the 0.5B model and 1152 for the 7B model in the first stage, and 768 for the 0.5B model and 480 for the 7B model in the second stage. The image tokenizer and original pre-trained LLM parameters were frozen in both stages, while other parameters were trainable. A consistent image resolution of 256x256 and the LDM-KL-16 [47] tokenizer were used throughout.
@@ 4.3 Comparison with Other Systems
Visual Understanding.
@@
In Table 1, we employ VLMEvalKit [12] to perform extensive evaluations on prevalent visual understanding benchmarks, encompassing a total of 18 such assessments (average score denoted by “AVE@18Und.” in Table 1, for details see Appendix 6.6.) including MMB [35], MME [15], POPE [31], SEED [26], MM-Vet [65], among others. Our method outperforms other joint image-text probabilistic models by a large margin, including Chameleon-7B, Show-o, VILA-U and our re-implemented version of Transfusion (denoted by “Transfusion*”). With merely 256 image token input, it achieves comparable or better performance compared to EMU3, which needs 4096 tokens per image. Even without using any pre-trained CLIP or diffusion models and with small resolution of 256 $\times$ 256, MMAR-7B presents comparable or better performance when compared to methods rely on pre-trained clip and diffusion models, including understanding-only methods like LLaVA-1.5, and non-probabilistic joint image understanding and generation models like EMU-2 and SEED-X.
Visual Generation.
We showcase the zero-shot FID [20] of the MMAR at the end of the second stage training, evaluated on the MSCOCO 30k dataset [32], in Table 3. It is worth noting that, even with only 3 epochs of training and using a less pure high-quality dataset due to the need to balance understanding performance, MMAR has been able to generate reasonable images, although it has not achieved very good generation effects. Furthermore, in order to verify that MMAR has the ability to generate high-quality images, we additionally train for 6 epochs on a pure high-quality image dataset as the third stage, where we replace Capfusion with JourneyDB [52] based on the second stage dataset. In the third stage, we only use image diffusion loss to further explore the model’s generative capabilities. We evaluate MMAR on FID@MJHQ-30K [28] and GenEval [17] benchmarks at the end of the third stage training, and show results in Table 4. It can be seen that after training with pure high-quality images, our model’s performance is discernibly on par with existing methods. Meanwhile, our image loss is still decreasing, which indicates that the image generation capability can still be improved. Notably, we also evaluated the visual understanding ability of the MMAR at the end of the third stage, and the overall average metric only decreased by 1%. For generated example images and visual understanding performance, please refer to Appendix 6.7 and 6.6.
Table 2: Ablation study on MMAR
[…]
MMAR-0.5B 48.45 882.1 34.56 36.6
w/ $\epsilon$-pred. 45.53 880.7 32.21 61.53
show-O like (w/VQ) 37.54 618.2 29.70 66.26
transfusion-like 29.47 594.3 28.26 95.38
4.4 Ablation Study
Table 2 demonstrates that switching to the more common $\epsilon$-prediction leads to a significant decrease in both visual understanding and generation quality, confirming the effectiveness of our optimal diffusion parameterization. To evaluate the efficacy of various image-text joint modeling techniques, we devise two distinct versions based on the MMAR framework: one employing LDM-VQ-16 tokenizer [47] for discrete token modeling (refer to show-O), and the other utilizing Transformer for diffusion modeling (refer to Transfusion). Implementation details are provided in the Appendix 6.1. Our test results are presented in Table 2. The transfusion-like version considerably underperforms the other two approaches in both understanding and generation aspects. This is attributed to the substantial loss of image information when jointly modeling image-text and operating with limited training epochs. In contrast, our full method consistently delivers superior results.